OpenStudy (anonymous):

Can someone help me? I'm not sure how to construct a polynomial function with three points, I want to learn how to do this and I want to solve the equation on my own, so if someone could explain with a random example.

OpenStudy (anonymous):

@ganeshie8 @ParthKohli @TuringTest @whpalmer4 @texaschic101 @eliassaab @chmvijay @charlotte123 @bookworm00981 @johnweldon1993 @rose21 @Ashleyisakitty

OpenStudy (anonymous):

Anyone... Please, help?

OpenStudy (anonymous):

Let the three points be (1,2),(-1,6) and (2,3) for example' Let \[ f(x) = a x^2 + b x +c \] be the polynomial. We need to find three unknowns a, b, c and we will be done. We need three equations in three unknowns. Can you write these three equations?

OpenStudy (anonymous):

\[ f(1)=a + b + c=1\\ f(-1)=a - b+c =6\\ f(2)= 4 a +2 b +c=3 \] Solve these equations and you are done. You find a=1, b=-2 and c=3